A simple pendulum has a period of 3.6 s. What is its length? Answer in meters.
1
Expert's answer
2011-02-24T05:16:56-0500
The period and the length of mathematical pendulum are related by the following expression:
T = 2 π √(L/g)
Thus the length is L = T2g/4π2 = 3.62* 9.8 / (4*3.142) = 3.22 [m]